a ballistic pendulum consists of a block of weed suspended by chords. when a bullet is fired into the block it is free to rise. how high does a 2kg block rise when 10gram bullet is traveling at 300 m/sec? show work
Answer #1

I am assuming block stick with weed and rises with it.

Use conservation of energy:

Initial kinetic energy = final potential energy

0.5*m1*v^2 = (m1+m2)*g*h

0.5*0.01*(300)^2= (0.01+2)*9.8*h

h=22.8 m

Answer: 22.8 m